Problem 1 Evaluate the line integral / x2 ds, where C is the line segment from (3,0) to (0,4) in the xy-plane.
Answer #1

Given line integral, 1 x ds, line segment(3,0) to,(0,4) From the given, r(t) =(1-t)(3,0)+t(0,4) r(t)=(3-3t, 4t) A parameteriz